# Do Vision–Language Models See Urban Scenes as People Do? An Urban Perception Benchmark

### v1

@@ Abstract
Understanding how people read city scenes can inform design and planning. We introduce a small benchmark for testing vision–language models (VLMs) on urban perception using 100 Montreal street images, evenly split between photographs and photorealistic synthetic scenes. Twelve participants from seven community groups supplied 230 annotation forms across 30 dimensions mixing physical attributes and subjective impressions. French responses were normalized to English. We evaluated seven VLMs in a zero-shot setup with a structured prompt and deterministic parser. We use accuracy for single-choice items and Jaccard overlap for multi-label items; human agreement uses Krippendorff’s $\alpha$ and pairwise Jaccard. Results suggest stronger model alignment on visible, objective properties than subjective appraisals. The top system (claude-sonnet) reaches macro 0.31 and mean Jaccard 0.48 on multi-label items. Higher human agreement coincides with better model scores. Synthetic images slightly lower scores. We release the benchmark, prompts, and harness for reproducible, uncertainty-aware evaluation in participatory urban analysis. [^1] The dataset will be released together with the images and the evaluation grids from participants.
@@ 1 Introduction
Recent vision–language models (VLMs) integrate image understanding with natural language reasoning (Achiam et al., 2023; Li et al., 2023; Dai et al., 2023; LLaVA Team, 2024; Pichai et al., 2023). They promise general, instruction-following analysis of images without task-specific training. However, there is limited evidence about how well such systems emulate human perception of urban scenes, particularly on subjective qualities. Existing evaluations of spatial knowledge in language models highlight both competence and brittleness (Gurnee and Tegmark, 2024; Roberts et al., 2023; Mai et al., 2023; Bhandari et al., 2023; Manvi et al., 2024). Bringing these strands together requires human-centered protocols that respect subjectivity, document uncertainty, and make evaluation reproducible (Mushkani et al., 2025f).
@@
  • RQ1 To what extent do contemporary VLMs align with human judgments on objective (visually grounded) versus subjective (appraisal) dimensions?
@@ 2 Related Work
Pairwise comparisons facilitated large-scale mapping of perceived urban qualities, starting with the “collaborative image of the city” (Salesses et al., 2013) and extended by Place Pulse 2.0 (Dubey et al., 2016; Mushkani et al., 2025c). Street-view imagery has subsequently supported tasks from estimating socioeconomic status to tracking urban change (Fan et al., 2023). At regional scales, multispectral imagery and machine learning have been used to reveal socioeconomic gradients (Jean et al., 2016; Yeh et al., 2020; Tatem, 2017). This body of work suggests that images carry strong signals of the built environment and social life while underscoring that perception is not reducible to a single ground truth (Mushkani and Koseki, 2025b).
@@ 3 Dataset, Participants, and Annotation Protocol
Figure 1: Self-reported participant context (counts). These categories were optional and are reported only in aggregate to characterize the annotator pool. Categories are not mutually exclusive and reflect intersectional identities; participants could select multiple identity markers, so counts exceed the number of participants Mushkani (2025b).
@@
The annotation schema comprises 30 dimensions spanning four families: physical setting (e.g., space typology, spatial configuration, lighting, vegetation, maintenance, signage, barriers), human presence and activity (e.g., human presence, types of activities, economic activities, accessibility features, visibility), built form and aesthetics (e.g., built environment, architectural style, aesthetic elements, cultural elements), and subjective impressions (e.g., overall impression). Each dimension was defined with concise descriptions and canonical label sets, distinguishing single-choice from multi-label items. To avoid identity inference, we use observational descriptors such as Presence of mobility aids rather than attributes like ethnicity or gender (Mushkani and Koseki, 2025a).
@@ 7 Discussion
The benchmark suggests a clear stratification by observability. Dimensions grounded in readily visible structure, such as spatial configuration or presence of vegetation, are detected more consistently even in a zero-shot setting. By contrast, low-scoring dimensions combine two challenges: the required evidence is either rare in our panels (Sustainability, Public Amenities) or it involves diffuse cultural and affective interpretation (Cultural Elements, Overall Impression). These results echo observations in urban studies that subjective impressions are plural and context dependent, and they support presenting distributions and uncertainty rather than single labels (Mushkani et al., 2025d).
@@ 8 Ethics and Privacy
This study was reviewed by the authors’ institutional ethics board and determined to be minimal risk (Mushkani, 2025c). All participants provided informed consent and were compensated. Annotations are released only in aggregate form with no personal data. Photographs were captured by the authors in public spaces; faces and license plates were automatically blurred and manually verified prior to use. Synthetic images were generated with Stable Diffusion XL.
@@ 9 Limitations
The dataset is intentionally small to enable careful documentation, community partnership, and reliability analysis. This limits statistical power for subgroup analyses and restricts the diversity of Montreal contexts. Annotations reflect a specific group of community members; other groups could disagree Mushkani (2025a). We use a deterministic parser for French–English normalization and CSV extraction; although validated, any remaining mapping errors would underrate models. Our evaluation is strictly zero-shot. Light-weight tuning with local data or multi-turn interaction may yield higher agreement but would assess a different capability. Finally, while we avoid personally identifying information in images and release only de-identified annotations, any work that links urban perception with place requires caution to avoid stigmatizing neighborhoods or groups.
@@ 10 Implications for practice and research
For practice, present-day VLMs may assist with factual components of streetscape audits, such as identifying seating, vegetation, and typology, and can pre-annotate images for human review. They are less dependable for subjective appraisals. Tools should therefore surface both model predictions and human reliability for each dimension, display uncertainty, and allow community members to contest labels (Mushkani et al., 2025a).
For research, the benchmark suggests two directions. The first is methodological: evaluation should normalize for inter-annotator variability, report both accuracy and distributional fit, and separate subjective from objective items. The second is technical: future models could incorporate structured visual geometry for spatial tasks and learn to express calibrated uncertainty for subjective appraisals. Participatory co-design with local organizations can help align model outputs with situated values (Mushkani et al., 2025b, e).
@@ 11 Conclusion
We introduced a community-grounded benchmark for evaluating vision–language models on urban perception using Montreal street-level scenes. The analysis shows stronger alignment on objective dimensions, weaker alignment on subjective appraisals, and a positive relationship between human reliability and model performance; we also observe a modest but consistent performance drop on photorealistic synthetic scenes. Taken together, these results provide baseline yardsticks for seven contemporary VLMs under a strict zero-shot protocol and a documented pipeline that emphasizes determinism and transparency. [^3] The dataset will be released together with the images and the evaluation grids from participants.
@@
We release prompts, normalization mappings, and a deterministic harness to support reproducible comparisons and community extension. Our aim is a living benchmark that helps the urban AI community track progress on human–AI alignment, encourages uncertainty-aware reporting, and keeps participatory values at the center of street-scale analysis. [^4] The dataset will be released together with the images and the evaluation grids from participants.
@@ Appendix A Perception grid
Table 2: Perception grid: dimensions, allowed values, and response type.
@@
Size (visual estimate) Small (<500 m²); Medium (500–2000 m²); Large (>2000 m²); Not applicable single
@@ Appendix B Prompt and parsing details
The inference script provides a deterministic system prompt that: (1) lists the 30 columns in the exact order shown in Table LABEL:tab:perception-grid; (2) requires only a single CSV line with 30 fields and no header or commentary; (3) instructs models to use Not applicable when unclear; and (4) specifies the semicolon rule for multi‑label items. Parsing is rule‑based: replies are split on commas (with a CSV fallback), trimmed and mapped to the codebook tokens, and padded/truncated to the expected length. Any non‑conforming reply is captured in Comments and excluded from scoring.
@@ Appendix C Implementation notes for reproducibility
    embeds local images as base‑64 data URIs in the chat message (no remote fetches);
@@
    enforces the one‑line CSV contract and a deterministic post‑processor that writes Image_ID + 30 fields + Comments to disk.
@@
Panels p1p1p5p5 consist of photorealistic synthetic scenes and p6p6p10p10 are real photographs (50 images each). Figure 10 breaks down agreement by source. We observe modest, consistent decreases on synthetic images across models, but the gap is small compared with cross‑dimension variation and does not change the ranking.

### v2

@@ Abstract
Understanding how people read city scenes can inform design and planning. We introduce a small benchmark for testing vision–language models (VLMs) on urban perception using 100 Montreal street images, evenly split between photographs and photorealistic synthetic scenes. Twelve participants from seven community groups supplied 230 annotation forms across 30 dimensions mixing physical attributes and subjective impressions. French responses were normalized to English. We evaluated seven VLMs in a zero-shot setup with a structured prompt and deterministic parser. We use accuracy for single-choice items and Jaccard overlap for multi-label items; human agreement uses Krippendorff’s $\alpha$ and pairwise Jaccard. Results suggest stronger model alignment on visible, objective properties than subjective appraisals. The top system (Claude-sonnet) reaches macro 0.31 and mean Jaccard 0.48 on multi-label items. Higher human agreement coincides with better model scores. Synthetic images slightly lower scores. We release the benchmark, prompts, and harness for reproducible, uncertainty-aware evaluation in participatory urban analysis.
Dataset is available: https://huggingface.co/datasets/rsdmu/urban-perception-benchmark
@@ 1 Introduction
Recent vision–language models (VLMs) integrate image understanding with natural language reasoning (Achiam et al., 2023; Li et al., 2023; Dai et al., 2023; LLaVA Team, 2024; Pichai et al., 2023). They promise general, instruction-following analysis of images without task-specific training. However, there is limited evidence about how well such systems emulate human perception of urban scenes, particularly on subjective qualities. Existing evaluations of spatial knowledge in language models highlight both competence and brittleness (Gurnee and Tegmark, 2024; Roberts et al., 2023; Mai et al., 2023; Bhandari et al., 2023; Manvi et al., 2024). Bringing these strands together requires human-centered protocols that respect subjectivity, document uncertainty, and make evaluation reproducible.
@@
  • RQ1 To what extent do VLMs align with human judgments on objective (visually grounded) versus subjective (appraisal) dimensions?
@@ 2 Related Work
Pairwise comparisons facilitated large-scale mapping of perceived urban qualities, starting with the “collaborative image of the city” (Salesses et al., 2013) and extended by Place Pulse 2.0 (Dubey et al., 2016). Street-view imagery has subsequently supported tasks from estimating socioeconomic status to tracking urban change (Fan et al., 2023). At regional scales, multispectral imagery and machine learning have been used to reveal socioeconomic gradients (Jean et al., 2016; Yeh et al., 2020; Tatem, 2017). This body of work suggests that images carry strong signals of the built environment and social life while underscoring that perception is not reducible to a single ground truth (Mushkani and Koseki, 2025).
@@ 3 Dataset, Participants, and Annotation Protocol
Figure 1: Self-reported participant context (counts). These categories were optional and are reported only in aggregate to characterize the annotator pool. Categories are not mutually exclusive and reflect intersectional identities; participants could select multiple identity markers, so counts exceed the number of participants.
@@
The annotation schema comprises 30 dimensions spanning four families: physical setting (e.g., space typology, spatial configuration, lighting, vegetation, maintenance, signage, barriers), human presence and activity (e.g., human presence, types of activities, economic activities, accessibility features, visibility), built form and aesthetics (e.g., built environment, architectural style, aesthetic elements, cultural elements), and subjective impressions (e.g., overall impression). Each dimension was defined with concise descriptions and canonical label sets, distinguishing single-choice from multi-label items. To avoid identity inference, we use observational descriptors such as Presence of mobility aids rather than attributes like ethnicity or gender.
@@ 7 Discussion
The benchmark suggests a clear stratification by observability. Dimensions grounded in readily visible structure, such as spatial configuration or presence of vegetation, are detected more consistently even in a zero-shot setting. By contrast, low-scoring dimensions combine two challenges: the required evidence is either rare in our panels (Sustainability, Public Amenities) or it involves diffuse cultural and affective interpretation (Cultural Elements, Overall Impression). These results echo observations in urban studies that subjective impressions are plural and context dependent, and they support presenting distributions and uncertainty rather than single labels (Mushkani et al., 2025).
@@ 8 Ethics and Privacy
All participants provided informed consent and were compensated. Annotations are released only in aggregate form with no personal data. Photographs were captured by the authors in public spaces; faces and license plates were automatically blurred and manually verified prior to use. Synthetic images were generated with Stable Diffusion XL.
@@ 9 Limitations
The dataset is intentionally small to enable careful documentation, community partnership, and reliability analysis. This limits statistical power for subgroup analyses and restricts the diversity of Montreal contexts. Annotations reflect a specific group of community members; other groups could disagree. We use a deterministic parser for French–English normalization and CSV extraction; although validated, any remaining mapping errors would underrate models. Our evaluation is strictly zero-shot. Light-weight tuning with local data or multi-turn interaction may yield higher agreement but would assess a different capability. Finally, while we avoid personally identifying information in images and release only de-identified annotations, any work that links urban perception with place requires caution to avoid stigmatizing neighborhoods or groups.
@@ 10 Implications for practice and research
For practice, present-day VLMs may assist with factual components of streetscape audits, such as identifying seating, vegetation, and typology, and can pre-annotate images for human review. They are less dependable for subjective appraisals. Tools should therefore surface both model predictions and human reliability for each dimension, display uncertainty, and allow community members to contest labels.
For research, the benchmark suggests two directions. The first is methodological: evaluation should normalize for inter-annotator variability, report both accuracy and distributional fit, and separate subjective from objective items. The second is technical: future models could incorporate structured visual geometry for spatial tasks and learn to express calibrated uncertainty for subjective appraisals. Participatory co-design with local organizations can help align model outputs with situated values.
@@ 11 Conclusion
We introduced a community-grounded benchmark for evaluating vision–language models on urban perception using Montreal street-level scenes. The analysis shows stronger alignment on objective dimensions, weaker alignment on subjective appraisals, and a positive relationship between human reliability and model performance; we also observe a modest but consistent performance drop on photorealistic synthetic scenes. Taken together, these results provide baseline yardsticks for seven contemporary VLMs under a strict zero-shot protocol and a documented pipeline that emphasizes determinism and transparency.
@@
We release prompts, normalization mappings, and a deterministic harness to support reproducible comparisons and community extension. Our aim is a living benchmark that helps the urban AI community track progress on human–AI alignment, encourages uncertainty-aware reporting, and keeps participatory values at the center of street-scale analysis.
@@ Appendix A Perception grid
Table 2: Perception grid: dimensions, allowed values, and type.
@@
Size (visual estimate) Small (<500 m2); Medium (500–2000 m2); Large (>2000 m2); Not applicable single
@@ Appendix B Prompt and parsing details
The inference script provides a deterministic system prompt that: (1) lists the 30 columns in the exact order shown in Table LABEL:tab:perception-grid; (2) requires only a single CSV line with 30 fields and no header or commentary; (3) instructs models to use Not applicable when unclear; and (4) specifies the semicolon rule for multi-label items. Parsing is rule-based: replies are split on commas (with a CSV fallback), trimmed and mapped to the codebook tokens, and padded/truncated to the expected length. Any non-conforming reply is captured in Comments and excluded from scoring.
@@ Appendix C Implementation notes for reproducibility
    embeds local images as base-64 data URIs in the chat message (no remote fetches);
@@
    enforces the one-line CSV contract and a deterministic post-processor that writes Image_ID + 30 fields + Comments to disk.
@@
Panels p1p1p5p5 consist of photorealistic synthetic scenes and p6p6p10p10 are real photographs (50 images each). Figure 10 breaks down agreement by source. We observe modest, consistent decreases on synthetic images across models, but the gap is small compared with cross-dimension variation and does not change the ranking.